OpenStudy (anonymous):

Solve for x: 3^x-1 = 27

OpenStudy (anonymous):

3^x-1 = 3^3 hence x-1=3 or x=1+3 or x=4

OpenStudy (anonymous):

What about ln x - ln 5 = 0

OpenStudy (anonymous):

ln x =ln 5 or x=5

OpenStudy (anonymous):

Solve the exponential eq. Algebraically. Round your result to three decimal places: 10^x = 570